# Double Yangians and quantum vertex algebras, I

### v1

@@ 5.1. Centrally extended double Yangian
$H_{i}^{-}(z)H_{j}^{+}(w)^{-1}=H_{j}^{+}(w)^{-1}H_{i}^{-}(z)\frac{(w-z-a_{j,i}\hbar-\kappa\hbar)(w-z+a_{j,i}\hbar+\kappa\hbar)}{(w-z+a_{j,i}\hbar-\kappa\hbar)(w-z-a_{j,i}\hbar+\kappa\hbar)},$
@@
(5.20) $[\log H_{i}^{+}(z),X_{j}^{\pm}(w)]=\pm X_{j}^{\pm}(w)\log\left(\frac{z-w+a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}{z-w-a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}\right)\!.$
@@
$[h_{i,\mathcal{Y}}^{+}(z),X_{j}^{+}(w)]=\$$X_{j}^{+}(w)F(\partial_{z})\log\left(\frac{z-w+a_{i,j}\hbar+\kappa\hbar}{z-w-a_{i,j}+\kappa\hbar}\right)$
@@
$[h_{i,\mathcal{Y}}^{+}(z),X_{j}^{-}(w)]=\$$-X_{j}^{-}(w)F(\partial_{z})\log\left(\frac{z-w+a_{i,j}\hbar}{z-w-a_{i,j}\hbar}\right)$
@@
(5.21) $[\log H_{i}^{-}(z),X_{j}^{\pm}(w)]=\pm X_{j}^{\pm}(w)\log\left(\frac{w-z-a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}{w-z+a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}\right)\!.$
@@
$[h_{i,\mathcal{Y}}^{-}(z),X_{j}^{+}(w)]=\$$-X_{j}^{+}(w)F(\partial_{z})\log\left(\frac{w-z-a_{i,j}\hbar+\kappa\hbar}{w-z+a_{i,j}\hbar+\kappa\hbar}\right)$
@@
$[h_{i,\mathcal{Y}}^{-}(z),X_{j}^{-}(w)]=\$$X_{j}^{-}(w)F(\partial_{z})\log\left(\frac{w-z-a_{i,j}\hbar}{w-z+a_{i,j}\hbar}\right)$
@@
$\ =X_{j}^{+}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w+\hbar\kappa)^{-1}+(w-z+\hbar\kappa)^{-1})$
$\ =x_{j,\mathcal{Y}}^{+}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w+\hbar\kappa)^{-1}+(w-z+\hbar\kappa)^{-1}),$
(5.23) $[h_{i,\mathcal{Y}}(z),X_{j}^{-}(w)]=-X_{j}^{-}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w)^{-1}+(w-z)^{-1}).$
@@
(5.25) $[h_{i,\mathcal{Y}}(z),x_{j,\mathcal{Y}}^{-}(w)]=-x_{j,\mathcal{Y}}^{-}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w+\kappa\hbar)^{-1}+(w-z+\kappa\hbar)^{-1}).$
For $r\in I$, setting
$\bar{X}_{r}(z)=X_{r}^{-}(z-\kappa\hbar),\ \ K_{r}(z)=H_{r}^{+}(z-\kappa\hbar/2)^{-1},$
we have $x_{r,\mathcal{Y}}^{-}(z)=\bar{X}_{r}(z)K_{r}(z).$ It is clear that (DY6) is equivalent to
(5.26) $(z-w+a_{r,s}\hbar)\bar{X}_{r}(z)\bar{X}_{s}(w)=(z-w-a_{r,s}\hbar)\bar{X}_{s}(w)\bar{X}_{r}(z)$
for $r,s\in I$. On the other hand, in the presence of (DY1) and (DY3) we have
(5.27) $K_{r}(z)K_{s}(w)=K_{s}(w)K_{r}(z),$
(5.28) $K_{r}(z)\bar{X}_{s}(w)=\left(\frac{z-w+a_{r,s}\hbar}{z-w-a_{r,s}\hbar}\right)\bar{X}_{s}(w)K_{r}(z).$
@@
(5.29) $(z-w-a_{r,s}\hbar)x_{r,\mathcal{Y}}^{-}(z)x_{s,\mathcal{Y}}^{-}(w)=(z-w+a_{r,s}\hbar)\bar{X}_{r}(z)\bar{X}_{s}(w)K_{r}(z)K_{s}(w),$
(5.30) $(z-w+a_{r,s}\hbar)x_{s,\mathcal{Y}}^{-}(w)x_{r,\mathcal{Y}}^{-}(z)=(z-w-a_{r,s}\hbar)\bar{X}_{s}(w)\bar{X}_{r}(z)K_{s}(w)K_{r}(z),$
noticing that $a_{r,s}=a_{s,r}$. Then it follows that (DY6) is equivalent to
(5.31) $(z-w-a_{r,s}\hbar)x_{r,\mathcal{Y}}^{-}(z)x_{s,\mathcal{Y}}^{-}(w)=(z-w+a_{r,s}\hbar)x_{s,\mathcal{Y}}^{-}(w)x_{r,\mathcal{Y}}^{-}(z).$
Now, we discuss (DY7). It is clear that (DY7) holds if and only if
(5.32) $\bar{X}_{i}(z)\bar{X}_{j}(w)=\bar{X}_{j}(w)\bar{X}_{i}(z)\quad\mbox{if }a_{i,j}=0.$
As $K_{r}(z)\in\mathcal{A}[[\hbar]][[z^{-1}]]$, (5.28) is equivalent to
(5.33) $\left(\frac{z-w-a_{r,s}\hbar}{z-w+a_{r,s}\hbar}\right)K_{r}(z)\bar{X}_{s}(w)=\bar{X}_{s}(w)K_{r}(z).$
Then using (5.28) we have
$x_{j,\mathcal{Y}}^{-}(w)x_{i,\mathcal{Y}}^{-}(z)$ $=\bar{X}_{j}(w)\bar{X}_{i}(z)K_{j}(w)K_{i}(z),$
$x_{i,\mathcal{Y}}^{-}(z)x_{j,\mathcal{Y}}^{-}(w)$ $=\bar{X}_{i}(z)K_{i}(z)\bar{X}_{j}(w)K_{j}(w)$
$=\bar{X}_{i}(z)\bar{X}_{j}(w)K_{i}(z)K_{j}(w),\quad\mbox{since }a_{i,j}=0.$
Thus, in the presence of (DY1) and (DY3), (5.32) is equivalent to
$x_{j,\mathcal{Y}}^{-}(w)x_{i,\mathcal{Y}}^{-}(z)=x_{i,\mathcal{Y}}^{-}(z)x_{j,\mathcal{Y}}^{-}(w)\quad\mbox{if }a_{i,j}=0.$
This gives the equivalence involving (DY7).
Next, we consider the Serre relations. As $x_{r,\mathcal{Y}}^{+}(z)=X_{r}^{+}(z)$ for $r\in I$, we only need to consider the Serre relation for $(x_{i,\mathcal{Y}}^{-}(z_{1}),x_{i,\mathcal{Y}}^{-}(z_{2}),x_{j,\mathcal{Y}}^{-}(w))$. With relation (5.26), by Lemma 4.6 we have
$(z-w)^{-1}(z-w+2\hbar)\bar{X}_{i}(z)\bar{X}_{i}(w)=(w-z)^{-1}(w-z+2\hbar)\bar{X}_{i}(w)\bar{X}_{i}(z).$
@@
${\mathopen{\overset{\bullet}{{{\bullet}}}}{\bar{X}_{i}(z_{1})\bar{X}_{i}(z_{2})\bar{X}_{j}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}={\mathopen{\overset{\bullet}{{{\bullet}}}}{\bar{X}_{i}(z_{2})\bar{X}_{i}(z_{1})\bar{X}_{j}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}.$
@@
$z_{1}^{-1}\delta\left(\frac{w+\hbar}{z_{1}}\right)z_{2}^{-1}\delta\left(\frac{w-\hbar}{z_{2}}\right){\mathopen{\overset{\bullet}{{{\bullet}}}}{x_{i,\mathcal{Y}}^{-}(z_{1})x_{i,\mathcal{Y}}^{-}(z_{2})x_{j,\mathcal{Y}}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
@@
Applying $\operatorname{Res}_{z_{1}}\operatorname{Res}_{z_{2}}$ we get
${\mathopen{\overset{\bullet}{{{\bullet}}}}{x_{i,\mathcal{Y}}^{-}(z_{1})x_{i,\mathcal{Y}}^{-}(z_{2})x_{j,\mathcal{Y}}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w+\hbar,z_{2}=w-\hbar}$
$=$ ${\mathopen{\overset{\bullet}{{{\bullet}}}}{X_{i}^{-}(z_{1})X_{i}^{-}(z_{2})X_{j}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w-\hbar,z_{2}=w+\hbar}K_{i}(w-\hbar)K_{i}(w+\hbar)K_{j}(w).$
From this we see that ${\mathopen{\overset{\bullet}{{{\bullet}}}}{x_{i,\mathcal{Y}}^{-}(z_{1})x_{i,\mathcal{Y}}^{-}(z_{2})x_{j,\mathcal{Y}}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w+\hbar,z_{2}=w-\hbar}=0$ if and only if
${\mathopen{\overset{\bullet}{{{\bullet}}}}{X_{i}^{-}(z_{1})X_{i}^{-}(z_{2})X_{j}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w-\hbar,z_{2}=w+\hbar}=0.$
Then by Proposition 4.8, the Serre relation for $(\bar{X}_{i}(z_{1}),\bar{X}_{i}(z_{2}),\bar{X}_{j}(w))$ holds if and only if the Serre relation for $(x_{i,\mathcal{Y}}^{-}(z_{1}),x_{i,\mathcal{Y}}^{-}(z_{2}),x_{j,\mathcal{Y}}^{-}(w))$ holds.
@@
$[X_{i}^{+}(z),\bar{X}_{j}(w)]=\delta_{i,j}\frac{1}{2\hbar}\left(\!K_{i}(w)^{-1}z^{-1}\delta\!\left(\frac{w}{z}\right)-H_{i}^{-}(w-3/2\kappa\hbar)z^{-1}\delta\!\left(\frac{w-2\kappa\hbar}{z}\right)\right)\!.$
@@
$X_{i}^{+}(z)K_{j}(w)=K_{j}(w)X_{i}^{+}(z)\left(\frac{w-z+a_{j,i}\hbar}{w-z-a_{j,i}\hbar}\right)\!.$
@@
$x_{i,\mathcal{Y}}^{+}(z)x_{j,\mathcal{Y}}^{-}(w)-\left(\frac{w-z+a_{j,i}\hbar}{w-z-a_{j,i}\hbar}\right)x_{j,\mathcal{Y}}^{-}(w)x_{i,\mathcal{Y}}^{+}(z)$
@@
$\delta_{i,j}\frac{1}{2\hbar}\left(\!z^{-1}\delta\!\left(\frac{w}{z}\right)-C_{i}(w)z^{-1}\delta\!\left(\frac{w-2\kappa\hbar}{z}\right)\!\right)\!$
@@
$=\$$[a_{i,j}]_{q^{\frac{\partial}{\partial{z_{2}}}}}[\kappa]_{q^{\frac{\partial}{\partial{z_{2}}}}}\left((z_{1}-z_{2}+\kappa\hbar)^{-2}-(z_{2}-z_{1}+\kappa\hbar)^{-2}\right)\!,$
@@
$=\$$\pm Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})[a_{i,j}]_{q^{\frac{\partial}{\partial{z_{2}}}}}\left((z_{1}-z_{2}+\kappa\hbar)^{-1}-(-z_{2}+z_{1}-\kappa\hbar)^{-1}\right)\!,$
$(z_{1}-z_{2}-a_{i,j}\hbar)Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1})Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})$
$=\$$(z_{1}-z_{2}+a_{i,j}\hbar)Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1})$
for $i,j\in I$ and
$Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1})Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})=Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1})$
if $a_{i,j}=0$. In case $a_{i,j}=-1$, by Corollary 4.11 the order-$2$ Serre relations hold for the triples $(Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1}),Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{2}),Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),w))$.
@@
$Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{-}(z),z_{2})-\!\left(\frac{z_{2}-z_{1}+a_{i,j}\hbar}{z_{2}-z_{1}-a_{i,j}\hbar}\right)\!Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{-}(z),z_{2})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})=0.$
@@
(5.38) $Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{-}(z),z_{2})-\!\left(\frac{z_{2}-z_{1}+2\hbar}{z_{2}-z_{1}-2\hbar}\right)\!Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{-}(z),z_{2})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})$
$=\$$\frac{1}{2\hbar}\!\left(\!z_{1}^{-1}\delta\!\left(\frac{z_{2}}{z_{1}}\right)-Y_{{\mathcal{E}}}\left(C_{i}(z),z_{2}\right)z_{1}^{-1}\delta\!\left(\frac{z_{2}-2\kappa\hbar}{z_{1}}\right)\!\right)\!.$
@@
$Y_{{\mathcal{E}}}^{-}(h_{i,\mathcal{Y}}(x),z),Y_{{\mathcal{E}}}^{+}(h_{i,\mathcal{Y}}(x),w)]=[2]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\kappa\hbar)^{-2}.$

### v2

@@ 5.1. Centrally extended double Yangian
$H_{i}^{-}(z)H_{j}^{+}(w)^{-1}=H_{j}^{+}(w)^{-1}H_{i}^{-}(z)\frac{(w-z-r_{j}a_{j,i}\hbar-\kappa\hbar)(w-z+r_{j}a_{j,i}\hbar+\kappa\hbar)}{(w-z+r_{j}a_{j,i}\hbar-\kappa\hbar)(w-z-r_{j}a_{j,i}\hbar+\kappa\hbar)},$
@@
(5.20) $[\log H_{i}^{+}(z),X_{j}^{\pm}(w)]=\pm X_{j}^{\pm}(w)\log\left(\frac{z-w+r_{i}a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}{z-w-r_{i}a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}\right)\!.$
@@
$[h_{i,\mathcal{Y}}^{+}(z),X_{j}^{+}(w)]=\$$X_{j}^{+}(w)F(\partial_{z})\log\left(\frac{z-w+r_{i}a_{i,j}\hbar+\kappa\hbar}{z-w-r_{i}a_{i,j}+\kappa\hbar}\right)$
@@
$[h_{i,\mathcal{Y}}^{+}(z),X_{j}^{-}(w)]=\$$-X_{j}^{-}(w)F(\partial_{z})\log\left(\frac{z-w+r_{i}a_{i,j}\hbar}{z-w-r_{i}a_{i,j}\hbar}\right)$
@@
(5.21) $[\log H_{i}^{-}(z),X_{j}^{\pm}(w)]=\pm X_{j}^{\pm}(w)\log\left(\frac{w-z-r_{i}a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}{w-z+r_{i}a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}\right)\!.$
@@
$[h_{i,\mathcal{Y}}^{-}(z),X_{j}^{+}(w)]=\$$-X_{j}^{+}(w)F(\partial_{z})\log\left(\frac{w-z-r_{i}a_{i,j}\hbar+\kappa\hbar}{w-z+r_{i}a_{i,j}\hbar+\kappa\hbar}\right)$
@@
$[h_{i,\mathcal{Y}}^{-}(z),X_{j}^{-}(w)]=\$$X_{j}^{-}(w)F(\partial_{z})\log\left(\frac{w-z-r_{i}a_{i,j}\hbar}{w-z+r_{i}a_{i,j}\hbar}\right)$
@@
$\ =X_{j}^{+}(w)[r_{i}a_{i,j}]_{q^{\partial_{w}}}((z-w+\hbar\kappa)^{-1}+(w-z+\hbar\kappa)^{-1})$
$\ =x_{j,\mathcal{Y}}^{+}(w)[r_{i}a_{i,j}]_{q^{\partial_{w}}}((z-w+\hbar\kappa)^{-1}+(w-z+\hbar\kappa)^{-1}),$
(5.23) $[h_{i,\mathcal{Y}}(z),X_{j}^{-}(w)]=-X_{j}^{-}(w)[r_{i}a_{i,j}]_{q^{\partial_{w}}}((z-w)^{-1}+(w-z)^{-1}).$
@@
(5.25) $[h_{i,\mathcal{Y}}(z),x_{j,\mathcal{Y}}^{-}(w)]=-x_{j,\mathcal{Y}}^{-}(w)[r_{i}a_{i,j}]_{q^{\partial_{w}}}((z-w+\kappa\hbar)^{-1}+(w-z+\kappa\hbar)^{-1}).$
For $s\in I$, setting
$\bar{X}_{s}(z)=X_{s}^{-}(z-\kappa\hbar),\ \ K_{s}(z)=H_{s}^{+}(z-\kappa\hbar/2)^{-1},$
we have $x_{s,\mathcal{Y}}^{-}(z)=\bar{X}_{s}(z)K_{s}(z).$ It is clear that (DY6) is equivalent to
(5.26) $(z-w+r_{s}a_{s,t}\hbar)\bar{X}_{s}(z)\bar{X}_{t}(w)=(z-w-r_{s}a_{s,t}\hbar)\bar{X}_{t}(w)\bar{X}_{s}(z)$
for $s,t\in I$. On the other hand, in the presence of (DY1) and (DY3) we have
(5.27) $K_{s}(z)K_{t}(w)=K_{t}(w)K_{s}(z),$
(5.28) $K_{s}(z)\bar{X}_{t}(w)=\left(\frac{z-w+r_{s}a_{s,t}\hbar}{z-w-r_{s}a_{s,t}\hbar}\right)\bar{X}_{t}(w)K_{s}(z).$
@@
(5.29) $(z-w-r_{s}a_{s,t}\hbar)x_{s,\mathcal{Y}}^{-}(z)x_{t,\mathcal{Y}}^{-}(w)=(z-w+r_{s}a_{s,t}\hbar)\bar{X}_{s}(z)\bar{X}_{t}(w)K_{s}(z)K_{t}(w),$
(5.30) $(z-w+r_{s}a_{s,t}\hbar)x_{t,\mathcal{Y}}^{-}(w)x_{s,\mathcal{Y}}^{-}(z)=(z-w-r_{s}a_{s,t}\hbar)\bar{X}_{t}(w)\bar{X}_{s}(z)K_{t}(w)K_{s}(z),$
noticing that $r_{s}a_{s,t}=r_{t}a_{t,s}$. Then it follows that (DY6) is equivalent to
(5.31) $(z-w-r_{s}a_{s,t}\hbar)x_{s,\mathcal{Y}}^{-}(z)x_{t,\mathcal{Y}}^{-}(w)=(z-w+r_{s}a_{s,t}\hbar)x_{t,\mathcal{Y}}^{-}(w)x_{s,\mathcal{Y}}^{-}(z).$
Next, we consider the Serre relations. As $x_{s,\mathcal{Y}}^{+}(z)=X_{s}^{+}(z)$ for $s\in I$, we only need to consider the Serre relation for $(x_{i,\mathcal{Y}}^{-}(z),x_{j,\mathcal{Y}}^{-}(w),1-a_{i,j})$. With relation (5.26), by Lemma 4.5 we have
$(z-w)^{-1}(z-w+2r_{i}\hbar)\bar{X}_{i}(z)\bar{X}_{i}(w)=(w-z)^{-1}(w-z+2r_{i}\hbar)\bar{X}_{i}(w)\bar{X}_{i}(z).$
@@
${\mathopen{\overset{\bullet}{{{\bullet}}}}{\bar{X}_{i}(z_{1})\bar{X}_{i}(z_{2})\cdots\bar{X}_{i}(z_{m_{i,j}})\bar{X}_{j}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}={\mathopen{\overset{\bullet}{{{\bullet}}}}{\bar{X}_{i}(z_{m_{i,j}})\cdots\bar{X}_{i}(z_{1})\bar{X}_{j}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}.$
@@
$\prod_{s=1}^{1-a_{i,j}}z_{s}^{-1}\delta\left(\frac{w-r_{i}(a_{i,j}+2s-2)\hbar}{z_{s}}\right){\mathopen{\overset{\bullet}{{{\bullet}}}}{x_{i,\mathcal{Y}}^{-}(z_{1})x_{i,\mathcal{Y}}^{-}(z_{2})\cdots x_{i,\mathcal{Y}}^{-}(z_{1-a_{i,j}})x_{j,\mathcal{Y}}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
@@
$\quad\times K_{i}(z_{1})\cdots K_{i}(z_{1-a_{i,j}})K_{j}(w).$
Applying $\operatorname{Res}_{z_{1}}\cdots\operatorname{Res}_{z_{1-a_{i,j}}}$ we get
${\mathopen{\overset{\bullet}{{{\bullet}}}}{x_{i}^{-}(w-r_{i}a_{i,j}\hbar)x_{i}^{-}(w-r_{i}(a_{i,j}+2)\hbar)\cdots x_{i}^{-}(w+r_{i}a_{i,j}\hbar)x_{j}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
$=$ ${\mathopen{\overset{\bullet}{{{\bullet}}}}{\bar{X}_{i}(w+r_{i}a_{i,j}\hbar)\bar{X}_{i}(w+r_{i}(a_{i,j}+2)\hbar)\cdots\bar{X}_{i}(w-r_{i}a_{i,j}\hbar)\bar{X}_{j}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
$\quad\times K_{i}(w-r_{i}a_{i,j}\hbar)K_{i}(w-r_{i}(a_{i,j}+2)\hbar)K_{i}(w+r_{i}a_{i,j}\hbar)K_{j}(w).$
From this we see that
${\mathopen{\overset{\bullet}{{{\bullet}}}}{x_{i}^{-}(w-r_{i}a_{i,j}\hbar)x_{i}^{-}(w-r_{i}(a_{i,j}+2)\hbar)\cdots x_{i}^{-}(w+r_{i}a_{i,j}\hbar)x_{j}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}=0$
if and only if
${\mathopen{\overset{\bullet}{{{\bullet}}}}{\bar{X}_{i}(w+r_{i}a_{i,j}\hbar)\bar{X}_{i}(w+r_{i}(a_{i,j}+2)\hbar)\cdots\bar{X}_{i}(w-r_{i}a_{i,j}\hbar)\bar{X}_{j}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}=0.$
Then by Proposition 4.13, the Serre-like relation for $(\bar{X}_{i}(z),\bar{X}_{j}(z),1-a_{i,j})$ holds if and only if the Serre-like relation for $(x_{i,\mathcal{Y}}^{-}(z),x_{j,\mathcal{Y}}^{-}(z),1-a_{i,j})$ holds.
@@
$[X_{i}^{+}(z),\bar{X}_{j}(w)]=\delta_{i,j}\frac{1}{2r_{i}\hbar}\left(\!K_{i}(w)^{-1}z^{-1}\delta\!\left(\frac{w}{z}\right)-H_{i}^{-}(w-3/2\kappa\hbar)z^{-1}\delta\!\left(\frac{w-2\kappa\hbar}{z}\right)\right)\!.$
@@
$X_{i}^{+}(z)K_{j}(w)=K_{j}(w)X_{i}^{+}(z)\left(\frac{w-z+r_{j}a_{j,i}\hbar}{w-z-r_{j}a_{j,i}\hbar}\right)\!.$
@@
$x_{i,\mathcal{Y}}^{+}(z)x_{j,\mathcal{Y}}^{-}(w)-\left(\frac{w-z+r_{j}a_{j,i}\hbar}{w-z-r_{j}a_{j,i}\hbar}\right)x_{j,\mathcal{Y}}^{-}(w)x_{i,\mathcal{Y}}^{+}(z)$
@@
$\delta_{i,j}\frac{1}{2r_{i}\hbar}\left(\!z^{-1}\delta\!\left(\frac{w}{z}\right)-C_{i}(w)z^{-1}\delta\!\left(\frac{w-2\kappa\hbar}{z}\right)\!\right)\!$
@@
$=\$$[r_{i}a_{i,j}]_{q^{\frac{\partial}{\partial{z_{2}}}}}[\kappa]_{q^{\frac{\partial}{\partial{z_{2}}}}}\left((z_{1}-z_{2}+\kappa\hbar)^{-2}-(z_{2}-z_{1}+\kappa\hbar)^{-2}\right)\!,$
@@
$=\$$\pm Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})[r_{i}a_{i,j}]_{q^{\frac{\partial}{\partial{z_{2}}}}}\left((z_{1}-z_{2}+\kappa\hbar)^{-1}-(-z_{2}+z_{1}-\kappa\hbar)^{-1}\right)\!,$
$(z_{1}-z_{2}-r_{i}a_{i,j}\hbar)Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1})Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})$
$=\$$(z_{1}-z_{2}+r_{i}a_{i,j}\hbar)Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1})$
for $i,j\in I$. In case $a_{i,j}\leq 0$, by Corollary 4.16 the Serre-like relations hold for the triples $(Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1}),Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{1}),1-a_{i,j})$.
@@
$Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{-}(z),z_{2})-\!\left(\frac{z_{2}-z_{1}+r_{i}a_{i,j}\hbar}{z_{2}-z_{1}-r_{i}a_{i,j}\hbar}\right)\!Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{-}(z),z_{2})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})=0.$
@@
(5.36) $Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{-}(z),z_{2})-\!\left(\frac{z_{2}-z_{1}+2r_{i}\hbar}{z_{2}-z_{1}-2r_{i}\hbar}\right)\!Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{-}(z),z_{2})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})$
$=\$$\frac{1}{2r_{i}\hbar}\!\left(\!z_{1}^{-1}\delta\!\left(\frac{z_{2}}{z_{1}}\right)-Y_{{\mathcal{E}}}\left(C_{i}(z),z_{2}\right)z_{1}^{-1}\delta\!\left(\frac{z_{2}-2\kappa\hbar}{z_{1}}\right)\!\right)\!.$
@@
$Y_{{\mathcal{E}}}^{-}(h_{i,\mathcal{Y}}(x),z),Y_{{\mathcal{E}}}^{+}(h_{i,\mathcal{Y}}(x),w)]=[2r_{i}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\kappa\hbar)^{-2}.$